OpenStudy (anonymous):

Solve the system of equations. 3x – 4y = 25 3x + 2y = 1 A. (1, –1) B. (3, –4) C. (–3, –9) D. (7, –1)

OpenStudy (solomonzelman):

subtract the second equation FROM the first

OpenStudy (anonymous):

WHICH ONE

OpenStudy (solomonzelman):

First - Second

OpenStudy (anonymous):

I came up with 1x + 2y = 24

OpenStudy (solomonzelman):

Subtract each thing (x, y, and constants - (numbers that aren't variables like 1,2,3....) ) separately. 3x – 4y = 25 - 3x + 2y = 1 -------------------------- -3x - 3x -4y - 2y = 25 - 1 0x -6y = 24 SO FAR SO GOOD?

OpenStudy (anonymous):

so now I solve it from here? 0x + -6y = 24

OpenStudy (solomonzelman):

yes. and 0x + -6y = 24 is just same as -6y = 24 1) Solve for y. 2) plug in the result for y into any of the ORIGINAL equations 3) Solve for x 4) write the answer as (x,y) where x is the x-value you get AND y, is the y-value you get.

OpenStudy (anonymous):

i came up with B (3,-4)